OpenStudy (anonymous):

in 1920, the record for a certain race was 45.3 sec. In 1960 it was 44.9 sec. let r (t) equal the record in the race and and (t)= the number of years since 1920 a find a linear function that fits the data r(t) = round to the nearest hundreth what is the predicted record for 2003 in sec what is it for 2006 in sec and in what year will the predicted record be 44.43 sec round to the nearest year.

OpenStudy (anonymous):

find the slope of the line as \[\frac{44.9-45.3}{40}=-\frac{.4}{40}=-.01\]]

OpenStudy (anonymous):

so you model says it goes down by .01 second per year after 1920

OpenStudy (andrew314):

y = -.01x + 45.3 2003; 44.47 sec 2006; 44.44 sec 2007 for 44.43 sec

OpenStudy (anonymous):

function would be \[r(t)=-.01t+45.3\] where t is years since 1920

OpenStudy (anonymous):

the rest is what andrew said!